OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6

OpenStudy (anonymous):

slope must be -1 : 3 = - 1/3 and to find intersection you substitute this slope and the point you R given 2 = -1/3 * 6 + B Here B is the intersection value we are looking for 2= -2 + B So this gives B = 4